# Elliptic and quadrangular flow of protons in the high baryon density region

### v1

@@ IV Summary
In summary, we have investigated the elliptic and quadrangular flow of protons in Au+Au collisions at $\sqrt{s_{NN}}$ = 2.4 GeV using the hadronic transport model SMASH. By incorporating the mean-field potential, our results qualitatively reproduce both the measured elliptic and quadrangular flow, as well as the ratio $v_{4}/v_{2}^{2}$ observed in the HADES experiment. This suggests that baryonic interactions may play a crucial role in shaping anisotropic flow in the high baryon density region.
@@
Finally, we explored the energy dependence of $v_{4}/v_{2}^{2}$ in the high baryon density region. Our results predict that the ratio remains close to 0.5 in this region and suggest that early-stage evolution in the high baryon density region resembles ideal fluid behavior. Nevertheless, a more rigorous investigation is warranted, both from theoretical and experimental perspectives, to further substantiate these results. These findings provide valuable insights for future experimental data from upcoming facilities such as FAIR, NICA, and HIAF.

### v2

@@ IV Summary
In summary, we have investigated the elliptic and quadrangular flow of protons in Au+Au collisions at $\sqrt{s_{NN}}$ = 2.4 GeV using the hadronic transport model SMASH. By incorporating the mean-field potential, our results qualitatively reproduce the measured elliptic and quadrangular flow, as well as the ratio $v_{4}/v_{2}^{2}$ observed in the HADES experiment. This suggests that baryonic interactions play a crucial role in shaping anisotropic flow in the high-baryon-density region.
@@
Lastly, we explored the energy dependence of $v_{4}/v_{2}^{2}$ in the high-baryon-density region. Our results suggest that the ratio remains constant and close to 0.5 in this regime, consistent with the observation from the HADES experiment. Further theoretical and experimental studies are needed to better understand the underlying mechanisms at low collision energies. These findings provide valuable insights into the anisotropic flow development at high baryon density and will help guide future experiments from upcoming facilities such as FAIR, NICA, and HIAF.